A 12 oz gatorade bottle costs $2.50 and a 8oz gatorade bottle costs $1.75. Which size bottle is the better buy?
My work: I divided 12/2.50 and got 4.8 (I'm guessing the is the number of ounces per dollar) and for the other one i did 8/$1.75 and got 4.57oz. Is this correct?
I think the seccond bottle is the better buy

Found 2 solutions by ikleyn, Theo:
Answer by ikleyn(52788)

.

Your calculations are correct, but the logic is false.


You have MORE liquid per dollar in the first case !



Answer by Theo(13342)

one 12 ounce bottle costs 2.50
one 8 ounce bottle costs 1.75.

what you did is possible, but can be confusing.

you got the number of ounces per dollar.

12 / 2.50 = 4.8 ounces per dollar.
8 / 1.75 = 4.57 ounces per dollar.

you then assumed that the less ounces per dollar was cheaper.
in fact, the more ounces per dollar is cheaper.

what you should have done is get the cost per ounce.

2.50 / 12 = .20833
1.75 / 8 = .21875

the 12 ounce can is cheaper because the cost per ounce is cheaper.

another way i do it sometimes is this.

12 ounce can is 2.50
8 ounce can is 1.75

i look for the least common multiple, which is 24.

24 = 2 * 12 = 2 * 2.50 = 5 dollars.
24 = 3 * 8 = 3 * 1.75 = 5.25 dollars.

12 ounce can is cheaper.

you had the right idea, but you should have gotten dollars / ounce, and not ounces / dollar.